# Tri-hypercharge versus tri-darkcharge

### v1

@@ I Introduction
such that each ordinary fermion generation $a$ is charged only under the corresponding $U(1)_{Y_{a}}$, similar to the decomposition of lepton number into generation lepton numbers and in the spirit of recent works [41, 42]. Since the right-handed neutrino singlets carry zero hypercharge, they remain uncharged under all $U(1)_{Y_{a}}$. To distinguish them, we introduce an additional dark gauge symmetry, $U(1)_{\mathsf{D}}$, under which the right-handed neutrinos are charged, while all SM fermions are neutral (i.e., $\mathsf{D}=0$), as suggested in [43, 28, 30].
@@
Second, assigning distinct dark charges $\delta_{1,2,3}$ to the right-handed neutrinos $\nu_{1,2,3R}$, anomaly cancellation for $[\mathrm{Gravity}]^{2}U(1)_{\mathsf{D}}$ and $U(1)_{\mathsf{D}}^{3}$ requires
@@
The remainder of this paper is organized as follows. In Section II, we present the fundamental aspects of the model, including the gauge symmetry, particle content, residual symmetries, Yukawa interactions, and scalar potential. Section III is devoted to the scalar and gauge boson mass spectra. The fermion mass spectrum is analyzed in Section IV. Constraints from electroweak precision observables, collider searches, and flavor physics are discussed in Sections V, VI, and VII, respectively. DM phenomenology is examined in Section VIII, and our conclusions are summarized in Section IX.
@@ II The fully flipped inert doublet model
As mentioned, this work considers a fully flipped inert doublet model where the scalar sector is enlarged by the inclusion of several gauge singlet scalars and an inert scalar doublet, while the fermionic content is added by charged vector like fermions and right-handed Majorana neutrinos, whose inclusion is necessary for the implementation of seesaw mechanisms that yields the small masses of the first and second generation of SM charged femions as well as tree level type I and one loop level radiative seesaw mechanisms that produces the tiny masses of the active neutrinos. The tree-level type I seesaw mechanism generates the atmospheric neutrino mass squared splittings, whereas the solar neutrino mass squared splittings arise from a radiative seesaw mechanism at one-loop level. Our theory is based on the gauge symmetry
@@
To break the extended gauge symmetry and generate the correct mass spectrum for the particles, we introduce several scalar fields in addition to the SM Higgs doublet $H=(H^{+}\,H^{0})^{T}$. First, a scalar singlet $\Phi$ is included to spontaneously break the $U(1)_{\mathsf{D}}$ symmetry down to the residual dark parity $P_{\mathsf{D}}$, while simultaneously generating Majorana masses for the right-handed neutrinos $\nu_{1,2R}$. Furthermore, four electrically neutral scalar singlets—$\varphi_{12}$, $\varphi_{23}$, $\phi_{12}$, and $\phi_{23}$—are introduced. These scalars, referred to as flavons, carry non-zero generation hypercharges (associated with $U(1)_{Y_{1,2,3}}$), arranged such that the total hypercharge remains zero. Their roles are twofold: (i) to break the generation hypercharge groups spontaneously down to the SM hypercharge $U(1)_{Y}$, and (ii) to explain the observed mass hierarchies and the small mixing angles in the quark sector. Each of these scalar fields acquires a vacuum expectation value (VEV), given by
@@
$H$ $({\bf 1,2})$ 00 00 $1/2$ 00 $+$
$\varphi_{23}$ $({\bf 1,1})$ 00 $1/6$ $-1/6$ 00 $+$
$\phi_{23}$ $({\bf 1,1})$ 00 $-1/2$ $1/2$ 00 $+$
$\varphi_{12}$ ($1({\bf 1,1})$/6161/61 / 6 $-1/6$ 00 00 $+$
$\phi_{12}$ ($1({\bf 1,1})$/212-1/2- 1 / 2 $1/2$ 00 00 $+$
$\Phi$ $({\bf 1,1})$ 00 00 00 $2$ $+$
$\eta$ $({\bf 1,2})$ 00 00 $-1/2$ $-1$-
$\rho$ $({\bf 1,1})$ 00 00 00 $1$-
Table 2: Scalar fields and their quantum numbers.
The spontaneous gauge symmetry breaking is implemented through the following way, $SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle v_{12}}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y_{1}+Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle v_{23}}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle\Lambda}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}\otimes R_{\mathsf{D}}\stackrel{{\scriptstyle v}}{{\longrightarrow}}SU(3)_{C}\otimes U(1)_{Q}\otimes R_{\mathsf{D}}$, in which we have assumed $\Lambda<v_{23}$ for the potential discovery of new physics at the LHC. Here $U(1)_{Q}$ is the electromagnetic symmetry and $R_{\mathsf{D}}$ is a residual symmetry of $U(1)_{\mathsf{D}}$ that conserves all the VEVs, $R_{\mathsf{D}}=(-1)^{k\mathsf{D}}$ for $k$ integer. This residual symmetry is automorphic to a discrete group, such as $R_{\mathsf{D}}\cong\mathbb{Z}_{2}=\{1,(-1)^{\mathsf{D}}\}$, for which $P_{\mathsf{D}}=(-1)^{\mathsf{D}}$ is called a residual dark parity of $U(1)_{\mathsf{D}}$. Under $P_{\mathsf{D}}$, the SM fields, $\nu_{3R}$, $\Phi$, $\varphi_{12,23}$, and $\phi_{12,23}$ transform trivially, i.e., $P_{\mathsf{D}}=1$, whereas $\nu_{1,2R}$, $\eta$, and $\rho$ transform nontrivially, i.e., $P_{\mathsf{D}}=-1$, as presented in Tables 1 and 2.
@@
$u_{12}$ $({\bf 3,1})$ $1/6$ $1/2$ 00 00 $+$
$u_{13}$ $({\bf 3,1})$ $1/6$ 00 $1/2$ 00 $+$
$u_{23}$ $({\bf 3,1})$ 00 $1/6$ $1/2$ 00 $+$
$d_{12}$ $({\bf 3,1})$ $1/6$ $-1/2$ 00 00 $+$
$d_{13}$ $({\bf 3,1})$ $1/6$ 00 $-1/2$ 00 $+$
$d_{23}$ $({\bf 3,1})$ 00 $1/6$ $-1/2$ 00 $+$
$e_{12}$ ($1({\bf 1,1})$/212-1/2- 1 / 2 $-1/2$ 00 00 $+$
$e_{13}$ ($1({\bf 1,1})$/212-1/2- 1 / 2 00 $-1/2$ 00 $+$
$e_{23}$ $({\bf 1,1})$ 00 $-1/2$ $-1/2$ 00 $+$
$\nu_{13}$ ($1({\bf 1,1})$/212-1/2- 1 / 2 00 $1/2$ 00 $+$
$\nu_{23}$ $({\bf 1,1})$ 00 $-1/2$ $1/2$ 00 $+$
Table 3: Vector-like fermions and their quantum numbers.
@@ III.1 Gauge sector
[…]
@@ III.2 Scalar sector
$2\mu_{6}^{2}+\lambda_{21}v^{2}+\lambda_{20}\Lambda^{2}+(\lambda_{18}+\lambda_{19})v_{23}^{2}+(2\lambda_{12}+\lambda_{13}+\lambda_{22})v_{12}^{2}$ $=$ $0$ (36)
@@ IV.2 Neutral fermion sector
With the aid of $M_{3},m_{\nu_{13,23}},v_{12,23}\gg v$ thus $M_{M}\gg M_{D}$, we can apply the seesaw formula to extract a $3\times 3$ mass submatrix for the light neutrino part, i.e., $M_{\nu}^{\mathrm{tree}}\simeq-M_{D}M_{M}^{-1}M_{D}^{T}$. Further, assuming $M_{3}\gg m_{\nu_{13}}\simeq m_{\nu_{23}}\simeq v_{12}$ and taking $z_{1}=z_{2}=1$, we find
@@
Notably, this texture can account for the large neutrino mixing angles while keeping all relevant Yukawa couplings of order $\mathcal{O}(1)$, as naturally expected in flavor models. It also reproduces the observed neutrino mass scale, $m_{\nu}\sim 0.1$ eV, with a heavy mass scale, $M_{3}\sim 10^{11}$ TeV. Importantly, this outcome does not require fine-tuning of small Yukawa couplings, nor does it demand the gauge symmetry-breaking scales $v_{12}$ and $v_{23}$ to be huge. Consequently, these scales can generally reside within the multi-TeV range.
It is worth emphasizing that the neutrino mass matrix $M_{\nu}^{\mathrm{tree}}$ generated at tree level via the seesaw mechanism has rank 1. As a result, it yields only one massive active neutrino and two massless ones, which is incompatible with current neutrino oscillation data [4]. Fortunately, the interactions of the third row of Eq. (13), together with the last two scalar couplings in Eq. (8), induce a one-loop radiative correction to the neutrino mass matrix, as illustrated in Fig. 2. In the mass eigenstate basis, the interactions relevant to this radiative contribution are given by:
@@
$M_{\nu}^{\mathrm{rad}}=\frac{(\kappa V_{1\alpha})^{2}M_{\alpha}}{32\pi^{2}}\left(\frac{c^{2}_{R}m^{2}_{R_{1}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{R_{1}}}}{M_{\alpha}^{2}-m^{2}_{R_{1}}}-\frac{c^{2}_{I}m^{2}_{I_{1}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{I_{1}}}}{M_{\alpha}^{2}-m^{2}_{I_{1}}}+\frac{s^{2}_{R}m^{2}_{R_{2}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{R_{2}}}}{M_{\alpha}^{2}-m^{2}_{R_{2}}}-\frac{s^{2}_{I}m^{2}_{I_{2}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{I_{2}}}}{M_{\alpha}^{2}-m^{2}_{I_{2}}}\right).$ (59)
@@
Finally, it is worth noting that the total neutrino mass matrix, given by $M_{\nu}^{\mathrm{tot}}=M_{\nu}^{\mathrm{tree}}+M_{\nu}^{\mathrm{rad}}$, has rank 2. This structure yields two massive active neutrinos and one massless neutrino, which is sufficient to accommodate current neutrino oscillation data [4]. This outcome is reminiscent of what one obtains in a minimal type-I seesaw/scotogenic/scotoseesaw extension of the SM with only two right-handed neutrinos [12, 45, 46, 19, 23]. The total neutrino mass matrix can be diagonalized via a unitary transformation: $V_{\nu_{L}}^{T}M_{\nu}^{\mathrm{tot}}V_{\nu_{L}}=\text{diag}(0,m_{2},m_{3})$, where $m_{2,3}$ are the physical neutrino masses, and $V_{\nu_{L}}$ is a unitary matrix that relates the gauge basis neutrino states $\nu_{aL}=(\nu_{1L},\nu_{2L},\nu_{3L})^{T}$ to the physical flavor eigenstates $\nu_{iL}=(\nu_{eL},\nu_{\mu L},\nu_{\tau L})^{T}$ via $\nu_{aL}=[V_{\nu_{L}}]_{ai}\nu_{iL}$. Consequently, the Pontecorvo–Maki–Nakagawa–Sakata (PMNS) matrix, which governs neutrino mixing in charged-current interactions, is given by $V_{\text{PMNS}}=V_{\nu_{L}}^{\dagger}V_{e_{L}}$, where $V_{e_{L}}$ is the unitary matrix that diagonalizes the charged lepton mass matrix.

### v2

@@ I Introduction
such that each ordinary fermion generation $a$ is charged only under the corresponding $U(1)_{Y_{a}}$, similar to the decomposition of lepton number into generation lepton numbers and in the spirit of recent works [41, 42, 43]. Since the right-handed neutrino singlets carry zero hypercharge, they remain uncharged under all $U(1)_{Y_{a}}$. To distinguish them, we introduce an additional dark gauge symmetry, $U(1)_{\mathsf{D}}$, under which the right-handed neutrinos are charged, while all SM fermions are neutral (i.e., $\mathsf{D}=0$), as suggested in [44, 28, 30].
@@
Second, assigning distinct dark charges $\delta_{1,2,3}$ to the right-handed neutrinos $\nu_{1,2,3R}$, anomaly cancellation for $\mathrm{Gravity}]^{2}U(1)_{\mathsf{D}}$ and $U(1)_{\mathsf{D}}^{3}$ requires
@@
We would like to emphasize that the main novelty of this work, compared with the recent studies [41, 42, 43], lies in simultaneously addressing the origin of neutrino masses and DM. Whereas in the models of Refs [41, 42] tiny active neutrino masses are generated through a type-I seesaw mechanism, in our model these masses arise from an interplay of tree-level type-I and one-loop radiative seesaw mechanism. Furthermore, our model provides an explanation for the hierarchy between the atmospheric and solar neutrino mass-squared splittings, since the former arises at tree level whereas the later originates at one loop. In our model, the radiative nature of the seesaw mechanism that yields the solar neutrino mass-squared difference is guaranteed by a residual parity $P_{\mathsf{D}}$ symmetry, resulting from the spontaneous breaking of the dark $U(1)_{\mathsf{D}}$ gauge symmetry. This residual parity $P_{\mathsf{D}}$ symmetry also ensures the stability of the DM candidates. In our framework, the successful implementation of the seesaw mechanisms that generate tiny active neutrino masses, together with the requirement of an anomaly-free $U(1)_{\mathsf{D}}$ dark gauge symmetry, naturally implies the existence of exactly three right-handed neutrinos and two vector-like neutral leptons. The parity-even and parity-odd neutral leptonic fields mediate a hybrid scotoseesaw mechanism that produces the tiny neutrino masses, while the lightest parity-odd neutral lepton serves as a DM candidate stabilized by the residual parity $P_{\mathsf{D}}$ symmetry. All SM fermions are neutral under $U(1)_{\mathsf{D}}$, whereas the three right-handed neutrinos carry charges $\mathsf{D}=1,-1,0$. As a result, the $U(1)_{\mathsf{D}}$ gauge boson does not couple directly to SM fermions. Distinctive collider signatures of the present model, compared with scenarios without $U(1)_{\mathsf{D}}$, thus arise mainly from the production of $P_{\mathsf{D}}$-odd right-handed neutrinos and scalars, leading to characteristic final states with leptons and missing energy.
The remainder of this paper is organized as follows. In Section II, we present the fundamental aspects of the model, including the gauge symmetry, particle content, residual symmetries, Yukawa interactions, and scalar potential. Section III is devoted to the scalar and gauge boson mass spectra, while the fermion mass spectrum is analyzed in Section IV. Constraints from electroweak precision observables, collider searches, and flavor physics are discussed in Sections V, VI, and VII, respectively. Section VIII examines dark matter phenomenology, and our conclusions are summarized in Section IX. Mathematical details concerning the general hypercharge decomposition are given in Appendix A. The vector and axial-vector couplings are discussed in Appendix B. Appendix C provides a detailed discussion of the one-loop Landau poles for the Abelian factors $U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}$.
@@ II The fully flipped inert doublet model
As mentioned, this work considers a fully flipped inert doublet model where the scalar sector is enlarged by the inclusion of several gauge singlet scalars and an inert scalar doublet, while the fermionic content is added by charged vector like fermions and right-handed Majorana neutrinos, whose inclusion is necessary for the implementation of seesaw mechanisms that yields the small masses of the first and second generation of SM charged femions as well as tree level type I and one loop level radiative seesaw mechanisms that produces the tiny masses of the active neutrinos. The tree-level type I seesaw mechanism generates the atmospheric neutrino mass squared splitting, whereas the solar neutrino mass squared difference arises from a radiative seesaw mechanism at one-loop level. Our theory is based on the gauge symmetry
@@
[…]
@@
$\langle\Phi\rangle$ $=$ $\frac{1}{\sqrt{2}}\Lambda,\hskip 14.22636pt\langle\varphi_{23}\rangle\approx\langle\phi_{23}\rangle\approx\frac{1}{\sqrt{2}}v_{23},\hskip 14.22636pt\langle\varphi_{12}\rangle\approx\langle\phi_{12}\rangle\approx\frac{1}{\sqrt{2}}v_{12},$ (6)
Field $SU(3)_{C}\otimes SU(2)_{L}$ $U(1)_{Y_{1}}$ $U(1)_{Y_{2}}$ $U(1)_{Y_{3}}$ $U(1)_{\mathsf{D}}$ $P_{\mathsf{D}}$ VEV [GeV] Roles/Purposes
$H$ $({\bf 1,2})$ 0 0 $1/2$ 0 ++ $v=246$
@@
$\varphi_{23}$ $({\bf 1,1})$ 0 $1/6$ 1/6-1/6 0 ++ $v_{23}\sim 10^{4}$
[…]
$\phi_{23}$ $({\bf 1,1})$ 0 1/2-1/2 $1/2$ 0 ++ $v_{23}\sim 10^{4}$
$\varphi_{12}$ $({\bf 1,1})$ $1/6$ 1/6-1/6 0 0 ++ $v_{12}\sim 10^{6}$
[…]
$\phi_{12}$ $({\bf 1,1})$ 1/2-1/2 $1/2$ 0 0 ++ $v_{12}\sim 10^{6}$
$\Phi$ $({\bf 1,1})$ 0 0 0 22 ++ $\Lambda\sim 10^{3}$
[…]
$\eta$ $({\bf 1,2})$ 0 0 1/2-1/2 1-1 - 0
[…]
$\rho$ $({\bf 1,1})$ 0 0 0 11 - 0
Table 2: Scalar fields, their quantum numbers, VEV, and roles/purposes in the model.
The spontaneous gauge symmetry breaking is implemented through the following way, $SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle v_{12}}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y_{12}\equiv Y_{1}+Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle v_{23}}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}\otimes U(1)_{\mathsf{D}}\stackrel{{\scriptstyle\Lambda}}{{\longrightarrow}}SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}\otimes R_{\mathsf{D}}\stackrel{{\scriptstyle v}}{{\longrightarrow}}SU(3)_{C}\otimes U(1)_{Q}\otimes R_{\mathsf{D}}$, in which we have assumed $\Lambda<v_{23}$ for the potential discovery of new physics at the LHC. Here $U(1)_{Q}$ is the electromagnetic symmetry and $R_{\mathsf{D}}$ is a residual symmetry of $U(1)_{\mathsf{D}}$ that conserves all the VEVs, $R_{\mathsf{D}}=(-1)^{k\mathsf{D}}$ for $k$ integer. This residual symmetry is automorphic to a discrete group, such as $R_{\mathsf{D}}\cong\mathbb{Z}_{2}=\{1,(-1)^{\mathsf{D}}\}$, for which $P_{\mathsf{D}}=(-1)^{\mathsf{D}}$ is called a residual dark parity of $U(1)_{\mathsf{D}}$. Under $P_{\mathsf{D}}$, the SM fields, $\nu_{3R}$, $\Phi$, $\varphi_{12,23}$, and $\phi_{12,23}$ transform trivially, i.e., $P_{\mathsf{D}}=1$, whereas $\nu_{1,2R}$, $\eta$, and $\rho$ transform nontrivially, i.e., $P_{\mathsf{D}}=-1$, as presented in Tables 1 and 2.
@@
Field $SU(3)_{C}\otimes SU(2)_{L}$ $U(1)_{Y_{1}}$ $U(1)_{Y_{2}}$ $U(1)_{Y_{3}}$ $U(1)_{\mathsf{D}}$ $P_{\mathsf{D}}$ Roles/Purposes
$u_{12}$ $({\bf 3,1})$ $1/6$ $1/2$ 0 0 ++
[…]
$u_{13}$ $({\bf 3,1})$ $1/6$ 0 $1/2$ 0 ++
$u_{23}$ $({\bf 3,1})$ 0 $1/6$ $1/2$ 0 ++
$d_{12}$ $({\bf 3,1})$ $1/6$ 1/2-1/2 0 0 ++
[…]
$d_{13}$ $({\bf 3,1})$ $1/6$ 0 1/2-1/2 0 ++
$d_{23}$ $({\bf 3,1})$ 0 $1/6$ 1/2-1/2 0 ++
$e_{12}$ $({\bf 1,1})$ 1/2-1/2 1/2-1/2 0 0 ++
[…]
$e_{13}$ $({\bf 1,1})$ 1/2-1/2 0 1/2-1/2 0 ++
$e_{23}$ $({\bf 1,1})$ 0 1/2-1/2 1/2-1/2 0 ++
$\nu_{13}$ $({\bf 1,1})$ 1/2-1/2 0 $1/2$ 0 ++
[…]
$\nu_{23}$ $({\bf 1,1})$ 0 1/2-1/2 $1/2$ 0 ++
Table 3: Vector-like fermions, their quantum numbers, and roles/purposes in the model.
@@ III.1 Gauge sector
$Z_{23\mu}=c_{\zeta}\mathcal{Z}_{1\mu}-s_{\zeta}\mathcal{Z}_{2\mu},\hskip 14.22636ptZ_{12\mu}=s_{\zeta}\mathcal{Z}_{1\mu}+c_{\zeta}\mathcal{Z}_{2\mu},$ (24)
@@ III.2 Scalar sector
$2\mu_{6}^{2}+\lambda_{21}v^{2}+\lambda_{20}\Lambda^{2}+(\lambda_{18}+\lambda_{19})v_{23}^{2}+(2\lambda_{12}+\lambda_{13}+\lambda_{22})v_{12}^{2}$ $=$ $0.$ (36)
@@
For completeness, we have analyzed the perturbativity of the extra Abelian gauge interactions in Appendix C, where the one-loop Landau poles of the four factors $U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}\otimes U(1)_{\mathsf{D}}$ are derived. The most restrictive case corresponds to $U(1)_{Y_{3}}$, where the corresponding Landau pole is found at the scale $\mu_{\text{LP},Y_{3}}\approx 10^{18}$ TeV, i.e. well above both the GUT and Planck scales. From this analysis we also extract upper bounds on the gauge couplings ensuring perturbativity up to given reference scales, namely $g_{3}\lesssim 0.846$ at the typical seesaw scale ($M_{3}\sim 10^{11}$ TeV), $g_{3}\lesssim 0.733$ at the GUT scale, and $g_{3}\lesssim 0.655$ at the Planck scale. These results justify the parameter ranges adopted in the phenomenological study of flavor observables, collider signatures, and DM in the following sections.
@@ IV Fermion mass and mixing
In this section, we study the generation of fermion masses and the structure of mixings. Since the model introduces several new energy scales, it is useful to begin with a global overview. Table 4 summarizes the characteristic scales and their typical values, highlighting the hierarchy that later appears in the effective operators. This overview provides a clear reference for the scale ratios relevant to the generation of active neutrino and SM charged fermion masses, as well as to the fermionic mixing pattern.
New scale(s) Typical value
$\Lambda,\,M_{1,2}$ $\mathcal{O}(1)\,\text{TeV}$
$v_{23}$ $\mathcal{O}(10)\,\text{TeV}$
$v_{12},\,m_{u_{23}},\,m_{\nu_{13,23}},\,m_{R_{1,2}},\,m_{I_{1,2}}$ $\mathcal{O}(10^{3})\,\text{TeV}$
$m_{u_{13}},\,m_{d_{13,23}},\,m_{e_{23}}$ $\mathcal{O}(10^{4})\,\text{TeV}$
$m_{u_{12}},\,m_{d_{12}},\,m_{e_{12,13}}$ $\mathcal{O}(10^{5})\,\text{TeV}$
$M_{3}$ $\mathcal{O}(10^{11})\,\text{TeV}$
Table 4: New energy scales appearing in the model together with their typical values.
@@ IV.2 Neutral fermion sector
With the aid of the hierarchy $M_{3},m_{\nu_{13,23}},v_{12,23}\gg v$, one finds $M_{M}\gg M_{D}$, so that the seesaw formula can be applied to extract a $3\times 3$ submatrix for the light neutrino sector. The resulting neutrino mass matrix takes a factorized form and can be written as the outer product of two vectors, i.e.,
$M_{\nu}^{\mathrm{tree}}$ $\simeq$ $-M_{D}M_{M}^{-1}M_{D}^{T}$ (56)
$=$ $\begin{pmatrix}-\frac{v^{2}v_{12}^{2}v_{23}^{2}x_{1}^{2}y_{1}^{2}z_{1}^{2}}{2F_{1}F_{2}}\\ \frac{m_{\nu_{13}}v^{2}v_{12}v_{23}^{2}x_{1}y_{1}y_{2}z_{1}^{2}}{2F_{1}F_{2}}\\ -\frac{v^{2}v_{12}v_{23}x_{1}y_{1}y_{3}z_{1}}{2F_{2}}\end{pmatrix}\begin{pmatrix}1&-\frac{m_{\nu_{13}}y_{2}}{v_{12}x_{1}y_{1}}&\frac{F_{1}y_{3}}{v_{12}v_{23}x_{1}y_{1}z_{1}}\end{pmatrix}$,
where $F_{1}=m_{\nu_{13}}m_{\nu_{23}}-x_{1}x_{2}v_{12}^{2}$ and $F_{2}=M_{3}F_{1}-\sqrt{2}m_{\nu_{13}}z_{1}z_{2}v_{23}^{2}$. Further, assuming $M_{3}\gg m_{\nu_{13}}\simeq m_{\nu_{23}}\simeq v_{12}$ and taking $z_{1}=z_{2}=1$, we obtain
@@
Notably, although most entries of $M_{\nu}^{\mathrm{tree}}$ are suppressed by the ratio $v_{23}/v_{12}\sim 10^{-2}$, this is compensated by the small denominator $(x_{1}x_{2}-1)\sim 10^{-2}$, so that all entries of the mass matrix can be of the same order. Consequently, the texture naturally yields large neutrino mixing angles without requiring fine-tuning of small Yukawa couplings. For representative values obtained in the numerical analysis below, $x_{1}\simeq 1.27$, $x_{2}\simeq 0.74$, $y_{1}\simeq 1.10$, $y_{2}\simeq 0.24$, $y_{3}\simeq 0.24$, the Yukawa couplings remain $\mathcal{O}(1)$, as typically expected in flavor models. Furthermore, the observed neutrino mass scale $m_{\nu}\sim 0.1$ eV is reproduced for a heavy mass scale $M_{3}\sim 10^{11}$ TeV, while the gauge symmetry-breaking scales $v_{12}$ and $v_{23}$ can lie in the multi-TeV range.
It is worth emphasizing that the tree-level neutrino mass matrix $M_{\nu}^{\mathrm{tree}}$ generated via the seesaw mechanism, takes the factorized form discussed above and is therefore manifestly of rank 1. Consequently, it predicts only one massive active neutrino, while the other two remain massless, in conflict with current neutrino oscillation data [4]. However, the interactions of the third row in Eq. (13), together with the last two scalar couplings in Eq. (8), induce a one-loop radiative correction to the $(3,3)$ entry of the neutrino mass matrix, as illustrated in Fig. 2. As a result, in our model the atmospheric neutrino mass-squared splitting is generated at tree level, whereas the solar neutrino mass-squared difference arises at one loop. In the mass eigenstate basis, the interactions relevant to this radiative contribution are given by:
@@
$M_{\nu}^{\mathrm{rad}}]_{33}=\frac{(\kappa V_{1\alpha})^{2}M_{\alpha}}{32\pi^{2}}\left(\frac{c^{2}_{R}m^{2}_{R_{1}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{R_{1}}}}{M_{\alpha}^{2}-m^{2}_{R_{1}}}-\frac{c^{2}_{I}m^{2}_{I_{1}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{I_{1}}}}{M_{\alpha}^{2}-m^{2}_{I_{1}}}+\frac{s^{2}_{R}m^{2}_{R_{2}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{R_{2}}}}{M_{\alpha}^{2}-m^{2}_{R_{2}}}-\frac{s^{2}_{I}m^{2}_{I_{2}}\ln\frac{M_{\alpha}^{2}}{m^{2}_{I_{2}}}}{M_{\alpha}^{2}-m^{2}_{I_{2}}}\right).$ (60)
@@
Finally, it is worth noting that the total neutrino mass matrix, given by
$M_{\nu}^{\mathrm{tot}}=M_{\nu}^{\mathrm{tree}}+M_{\nu}^{\mathrm{rad}},$ (61)
has rank 2. This structure yields two massive active neutrinos and one massless neutrino, which is sufficient to accommodate current neutrino oscillation data [4]. This outcome is reminiscent of what one obtains in a minimal type-I seesaw/scotogenic/scotoseesaw extension of the SM with only two right-handed neutrinos [12, 46, 47, 19, 23]. The total neutrino mass matrix can be diagonalized via a unitary transformation: $V_{\nu_{L}}^{T}M_{\nu}^{\mathrm{tot}}V_{\nu_{L}}=\text{diag}(0,m_{2},m_{3})$, where $m_{2,3}$ are the physical neutrino masses, and $V_{\nu_{L}}$ is a unitary matrix that relates the gauge basis neutrino states $\nu_{aL}=(\nu_{1L},\nu_{2L},\nu_{3L})^{T}$ to the physical flavor eigenstates $\nu_{iL}=(\nu_{eL},\nu_{\mu L},\nu_{\tau L})^{T}$ via $\nu_{aL}=[V_{\nu_{L}}]_{ai}\nu_{iL}$. Consequently, the Pontecorvo–Maki–Nakagawa–Sakata (PMNS) matrix, which governs neutrino mixing in charged-current interactions, is given by $V_{\text{PMNS}}=V_{\nu_{L}}^{\dagger}V_{e_{L}}$, where $V_{e_{L}}$ is the unitary matrix that diagonalizes the charged lepton mass matrix.